Please answer in detail on how these 2 answers are gotten. Include Heat formula (or any formula) and all math used to get to answer

Calculate the mass, in grams, for each of the following using the table below Table Specific Heats of Some Substances Specific Heat (J/g C) Substance Aliminium, Al(s) Copper, Cu(s) Gold, Au(s) ron, Fe(s) Silver. Ag(s) Titanium, Ti(s) 0.897 0.385 0.129 0.452 0.235 0.523 2.04 2.46 0.864 4.184 2.03 Elements Compounds Ammonia, NHs(g) Ethanol, C2H5OHl) Sodium chloride, NaCl(s) Water, H2Od) Water, H2O(s) Part A a sample of aluminum (Al) that absorbs 8.80 kJ when heated from 18.5 °C to 27.8°C Express your answer with the appropriate units 四 ? Value Units Submit Request Answer Part B a sample of titanium (Ti) that loses 1.49x104 J when it cools from 182 C to 42 C Express your answer with the appropriate units Value Units Submit Request Answer
Answer #1

part-A

q   = mc\Deltat

c   = 0.897J/g-c

q   = 8.80Kj   = 8800J

q   = mc\Deltat

8800 = m*0.897*(27.8-18.5)

m     = 8800/8.3421   = 1054.9g

mass of Al   = 1054.9g

part-B

q = mc\Deltat

q    = 1.49*10^4J   = 14900J

c   = 0.523J/g-c

q = mc\Deltat

-14900   = m*0.523*(42-182)

-14900   = m*0.523*-140

m    = -14900/0.523*-140   = 203.5g

mass of Ti    = 203.5g >>>>answer